# Failure Makes the Agent Stronger: Enhancing Accuracy through Structured Reflection for Reliable Tool Interactions

### v2

@@ 9.3 Core Reward with Similarity Backoff: Signal and Variance Control
Let $\mathcal{B}=\{R_{\text{core}}<\varepsilon\}$ with $\mathbb{P}(\mathcal{B})=p>0$. Assume $\mathrm{Sim}(\mathrm{concat}(C),\mathrm{concat}(G))\geq\sigma$ a.s. on $\mathcal{B}$ for some $\sigma>0$, and $\mathbb{E}\big[\|\nabla_{\theta}\log\pi_{\theta}(\cdot)\|^{2}\,\mathbf{1}_{\mathcal{B}}\big]>0$. Then
$\mathbb{E}\!\left[\ \big\|R_{\text{total}}\,\nabla_{\theta}\log\pi_{\theta}(\cdot)\big\|^{2}\ \right]\ \geq\ (w_{\text{b}}\sigma)^{2}\ \mathbb{E}\!\left[\ \big\|\nabla_{\theta}\log\pi_{\theta}(\cdot)\big\|^{2}\,\mathbf{1}_{\mathcal{B}}\ \right]\ >\ 0.$ (39)
Implication. When $R_{\text{core}}$ requently approaches 0 (in the early stages of training), the backoff branch ensures that the second moment of the gradients does not degenerate; combined with the variance upper bound from Lemma 3, this helps stabilize the optimization updates.
9.4 Sequence-Level Importance Sampling and Clipping
Let the sampled completion be $o=(o_{1},\dots,o_{T})$, and define the sequence-level (geometric-mean, length-normalized) ratio
$r_{\mathrm{seq}}(\theta)=\left(\prod_{t=1}^{T}\frac{\pi_{\theta}\!\left(o_{t}\mid q,\,o_{<t}\right)}{\pi_{\theta_{\text{old}}}\!\left(o_{t}\mid q,\,o_{<t}\right)}\right)^{\!1/T}\;=\;\exp\!\Bigl(\tfrac{1}{T}\sum_{t=1}^{T}\log\rho_{t}\Bigr),\quad\rho_{t}=\frac{\pi_{\theta}(o_{t}\mid\cdot)}{\pi_{\theta_{\text{old}}}(o_{t}\mid\cdot)}.$ (40)
@@ 9.4 Sequence-Level Importance Sampling and Clipping
9.5 Dynamic Filtering of Prompt Groups (DAPO-style)
@@
$\hat{A}_{i}\;=\;\frac{R_{i}-\bar{R}}{s_{R}},\qquad\bar{R}=\tfrac{1}{G}\sum_{j=1}^{G}R_{j},\ \ s_{R}=\sqrt{\tfrac{1}{G}\sum_{j=1}^{G}(R_{j}-\bar{R})^{2}}\;>\;0.$ (42)
@@
$\mathcal{S}\;=\;\big\{\,i:\ |\hat{A}_{i}|>\tau_{\text{adv}}\,\big\},\qquad 0<|\mathcal{S}|<G,\qquad\mathrm{Var}(\{R_{i}\}_{i=1}^{G})>\tau_{\text{var}}>0.$ (43)
@@
$\ell_{i}(\theta)\;=\;\min\Big\{\,r_{\mathrm{seq},i}(\theta)\,\hat{A}_{i},\;\mathrm{clip}\!\big(r_{\mathrm{seq},i}(\theta),1-\varepsilon_{\text{low}},1+\varepsilon_{\text{high}}\big)\,\hat{A}_{i}\,\Big\},$ (44)
@@
$\big\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\big\|\;\leq\;B_{\pi},\quad r_{\mathrm{seq},i}(\theta)\in[\,1-\varepsilon_{\text{low}},\,1+\varepsilon_{\text{high}}\,].$ ($\star$)
@@
Since $r_{\mathrm{seq},i}(\theta)$ is the geometric mean of token ratios,
$\nabla_{\theta}r_{\mathrm{seq},i}(\theta)\;=\;r_{\mathrm{seq},i}(\theta)\,\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}\mid q,o_{i,<t}).$ (45)
@@
(a) If $\hat{A}_{i}=0$, removing $o_{i}$ leaves the group-wise expected gradient unchanged.
(b) If $|\hat{A}_{i}|\leq\tau_{\text{adv}}$, then, for any $\theta$,
$\big\|\ \mathbb{E}\,[\,g_{i}(\theta)\,]\ \big\|\;\leq\;C_{\psi}\,\tau_{\text{adv}},\qquad\mathbb{E}\!\left[\ \|g_{i}(\theta)\|^{2}\ \right]\;\leq\;C_{\psi}^{2}\,\tau_{\text{adv}}^{\,2}.$ (47)
@@
$\tilde{g}(\theta)\;=\;\frac{1}{G}\sum_{i\in\mathcal{S}}g_{i}(\theta),\qquad g(\theta)\;=\;\frac{1}{G}\sum_{i=1}^{G}g_{i}(\theta)\ \ \text{(unfiltered)}.$ (48)
@@
9.6 Convergence Considerations for the Clipped Sequence-Level Objective
@@
Let $o=(o_{1},\dots,o_{|o|})$ and $r_{\mathrm{seq}}(\theta)$ denote the (clipped) sequence ratio. Then
$\nabla_{\theta}r_{\mathrm{seq}}(\theta)\;=\;r_{\mathrm{seq}}(\theta)\,\frac{1}{|o|}\sum_{t=1}^{|o|}\nabla_{\theta}\log\pi_{\theta}(o_{t}\!\mid\!q,o_{<t}),\qquad\big\|\nabla_{\theta}r_{\mathrm{seq}}(\theta)\big\|\leq(1+\varepsilon_{\text{high}})\,B_{\pi}.$ (52)

### v3

@@ A.8.3 Core Reward with Similarity Backoff: Signal and Variance Control
Let $\mathcal{B}:=\{R_{\text{core}}<\varepsilon\}$ with $\mathbb{P}(\mathcal{B})=p>0$. Let $g_{\theta}:=\nabla_{\theta}\log\pi_{\theta}(\cdot)$. Assume $S_{\text{b}}\geq\sigma$ a.s. on $\mathcal{B}$ for some $\sigma>0$ and $\mathbb{E}\!\left[\|g_{\theta}\|^{2}\,\mathbf{1}_{\mathcal{B}}\right]>0$. Then
$\mathbb{E}\!\left[\ \big\|R_{\text{total}}\,g_{\theta}\big\|^{2}\ \right]$ $\geq(w_{\text{b}}\sigma)^{2}\ \mathbb{E}\!\left[\ \|g_{\theta}\|^{2}\,\mathbf{1}_{\mathcal{B}}\ \right]$ (39)
$>0.$
Implication. When $R_{\text{core}}$ frequently approaches 0 (early in training), the backoff branch prevents the gradient second moment from degenerating; together with the variance upper bound in Lemma 3, this stabilizes the optimization updates.
A.8.4 Sequence-Level Importance Sampling and Clipping
Let the sampled completion be $o=(o_{1},\dots,o_{T})$, and define the sequence-level (geometric-mean, length-normalized) ratio:
$\begin{split}r_{\mathrm{seq}}(\theta)&=\left(\prod_{t=1}^{T}\frac{\pi_{\theta}\!\left(o_{t}\mid q,\,o_{<t}\right)}{\pi_{\theta_{\text{old}}}\!\left(o_{t}\mid q,\,o_{<t}\right)}\right)^{\!1/T}\\ &=\exp\!\left(\frac{1}{T}\sum_{t=1}^{T}\log\rho_{t}\right),\\ \rho_{t}&:=\frac{\pi_{\theta}\!\left(o_{t}\mid q,\,o_{<t}\right)}{\pi_{\theta_{\text{old}}}\!\left(o_{t}\mid q,\,o_{<t}\right)}.\end{split}$ (40)
@@ A.8.4 Sequence-Level Importance Sampling and Clipping
A.8.5 Dynamic Filtering of Prompt Groups (DAPO-style)
@@
$\hat{A}_{i}$ $=\frac{R_{i}-\bar{R}}{s_{R}},$ (42)
$\bar{R}$ $=\frac{1}{G}\sum_{j=1}^{G}R_{j},$
$s_{R}$ $=\sqrt{\frac{1}{G}\sum_{j=1}^{G}(R_{j}-\bar{R})^{2}}\;>0.$
@@
$\mathcal{S}$ $=\big\{\,i:|\hat{A}_{i}|>\tau_{\text{adv}}\,\big\},$ (43)
$0<|\mathcal{S}|$ $<G,$
$\mathrm{Var}\!\big(\{R_{i}\}_{i=1}^{G}\big)$ $>\tau_{\text{var}}\;>0.$
@@
$\ell_{i}(\theta)$ $=\min\!\Big(r_{\mathrm{seq},i}(\theta),\,\bar{r}_{\mathrm{seq},i}(\theta)\Big)\,\hat{A}_{i},$ (44)
$\bar{r}_{\mathrm{seq},i}(\theta)$ $=\mathrm{clip}\!\Big(r_{\mathrm{seq},i}(\theta),1-\varepsilon_{\text{low}},1+\varepsilon_{\text{high}}\Big).$
@@
$\big\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\big\|$ $\leq B_{\pi},$ ($\star$)
$r_{\mathrm{seq},i}(\theta)$ [ $\in\big[1-\varepsilon_{\text{low}},1+\varepsilon_{\text{high}}\,\big].$
@@
Since $r_{\mathrm{seq},i}(\theta)$ is the geometric mean of token ratios, let $T_{i}:=|o_{i}|$ and $h_{i,t}:=(q,o_{i,<t})$. Then
$\nabla_{\theta}r_{\mathrm{seq},i}(\theta)$ $=r_{\mathrm{seq},i}(\theta)\,\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}\mid h_{i,t})$ (45)
$=\frac{r_{\mathrm{seq},i}(\theta)}{T_{i}}\sum_{t=1}^{T_{i}}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}\mid h_{i,t}).$
@@
  1. (a)
    If $\hat{A}_{i}=0$, removing $o_{i}$ leaves the group-wise expected gradient unchanged.
  2. (b)
    If $|\hat{A}_{i}|\leq\tau_{\text{adv}}$, then for any $\theta$,
    $\big\|\,\mathbb{E}[\,g_{i}(\theta)\,]\,\big\|$ $\leq C_{\psi}\,\tau_{\text{adv}},$ (47)
    $\mathbb{E}\!\left[\ \|g_{i}(\theta)\|^{2}\ \right]$ $\leq C_{\psi}^{2}\,\tau_{\text{adv}}^{\,2}.$
@@
$\tilde{g}(\theta)$ $=\frac{1}{G}\sum_{i\in\mathcal{S}}g_{i}(\theta),$ (48)
$g(\theta)$ $=\frac{1}{G}\sum_{i=1}^{G}g_{i}(\theta)\qquad\text{(unfiltered)}.$
@@
A.8.6 Convergence Considerations for the Clipped Sequence-Level Objective
@@
Let $o=(o_{1},\dots,o_{T})$ with $T:=|o|$, and let $h_{t}:=(q,o_{<t})$. Let $r_{\mathrm{seq}}(\theta)$ denote the (clipped) sequence ratio. Then
$\nabla_{\theta}r_{\mathrm{seq}}(\theta)$ $=\frac{r_{\mathrm{seq}}(\theta)}{T}\sum_{t=1}^{T}\nabla_{\theta}\log\pi_{\theta}(o_{t}\mid h_{t}),$ (52)
$\big\|\nabla_{\theta}r_{\mathrm{seq}}(\theta)\big\|$ $\leq(1+\varepsilon_{\text{high}})\,B_{\pi}.$